# Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models

### v1

@@
Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models
Parham Rezaei , Farzan Farnia , Cheuk Ting Li Department of Computer Engineering, Sharif University of Technology, parham.rezaei@sharif.eduDepartment of Computer Science and Engineering, The Chinese University of Hong Kong, farnia@cse.cuhk.edu.hkDepartment of Information Engineering, The Chinese University of Hong Kong, ctli@ie.cuhk.edu.hk
The availability of multiple training algorithms and architectures for generative models requires a selection mechanism to form a single model over a group of well-trained generation models. The selection task is commonly addressed by identifying the model that maximizes an evaluation score based on the diversity and quality of the generated data. However, such a best-model identification approach overlooks the possibility that a mixture of available models can outperform each individual model. In this work, we explore the selection of a mixture of multiple generative models and formulate a quadratic optimization problem to find an optimal mixture model achieving the maximum of kernel-based evaluation scores including kernel inception distance (KID) and Rényi kernel entropy (RKE). To identify the optimal mixture of the models using the fewest possible sample queries, we propose an online learning approach called Mixture Upper Confidence Bound (Mixture-UCB). Specifically, our proposed online learning method can be extended to every convex quadratic function of the mixture weights, for which we prove a concentration bound to enable the application of the UCB approach. We prove a regret bound for the proposed Mixture-UCB algorithm and perform several numerical experiments to show the success of the proposed Mixture-UCB method in finding the optimal mixture of text-based and image-based generative models. The codebase is available at https://github.com/rezaei-parham/mixture-ucb.
@@ 1 Introduction
The rapid advancements in generative modeling have created a need for mechanisms to combine multiple well-trained generative models, each developed using different algorithms and architectures, into a single unified model. A common approach for creating such a unified model is to evaluate assessment scores that quantify the diversity and quality of the generated data and then select the model with the highest score. This best-model identification strategy has been widely adopted for the selection of generative models across various domains, including image, text, audio, and video generation.
Figure 1: Visual comparison of the diversity across individual arms and the optimal mixture for images generated using models Kandinsky 3, Stable Diffusion 3-medium, and PixArt-$\alpha$ with the prompt ”Red bird, cartoon style”
Existing model selection frameworks typically perform an offline selection, where they have access to a sufficiently large number of samples from each generative model and estimate the evaluation score based on these samples. However, in many practical scenarios, generating a large sample set from sub-optimal models can be computationally costly, especially if the evaluator can identify their lack of optimality using fewer samples. In such cases, the evaluator can adopt an online learning approach and frame the problem as a multi-armed bandit (MAB) task. An existing approach is successive halving [1, 2, 3],111[2] focused on applying successive halving on hyperparameter optimization for supervised learning, whereas [3] focused on generative models using maximum mean discrepancy [4] as the score. where the models are evaluated using a fixed budget, the worst half of the models are removed, and we repeat the process until one model is left. Also, the recent work by Hu et al. [5] attempts to solve the online model selection problem using an upper confidence bound (UCB) method to identify the generative model with the highest evaluation score. The numerical results in this work indicate the effectiveness of MAB algorithms in reducing sample generation costs from sub-optimal models.
@@
On the other hand, when the model selection task is handled by an online learning algorithm, the algorithm may choose different models at different iterations, resulting in generated data that follow a mixture of the distributions of these generative models. Note that in a standard MAB algorithm, the goal is to eventually converge to a single arm. Successive halving [1, 2, 3] and the standard UCB algorithm adopted by [5] will ultimately select only one generative model after a sufficient number of iterations. However, the diversity scores of the generated data could be higher when the sample generation follows a mixture of models rather than a single model. This observation leads to the following question: could the diversity of generated data be improved by applying MAB algorithms to multiple generative models, if the algorithm aims to find the best mixture of the models?
@@
In this work, we aim to address the above question by finding the optimal mixture $\sum_{i=1}^{m}\alpha_{i}P_{i}$ of $m$ generative models with distributions $P_{1},\ldots,P_{m}$, which would produce a higher evaluation score compared to each individual model. Specifically, we focus on addressing this task in an online learning setting, where we pick a model to generate a sample at each round. To address this problem and develop an MAB algorithm to find the best mixture model, we concentrate on evaluation scores that are quadratic functions of the generated data. As we show in this work, formulating the optimization problem for a quadratic score function results in a quadratic online convex optimization problem that can be efficiently solved using the online gradient descent algorithm. More importantly, we establish a concentration bound for the quadratic function of the mixture weights, which enables us to extend the UCB algorithm for the online selection of the mixture weights.
@@
Specifically, we focus on evaluation scores that reduce to a quadratic function of the generative model’s distribution, including the kernel-based Maximum Mean Discrepancy (MMD) [4], Kernel Inception Distance (KID) [6] and Rényi Kernel Entropy (RKE) [7] scores, as well as the quality-measuring Precision [8, 9] and Density [10] scores, which are linear functions of the generative distribution. Among these scores, RKE provides a reference-free entropy function for assessing the diversity of generated data, making it suitable for quantifying the variety of generated samples. Our mixture-based online learning framework can therefore be applied to find the mixture model with the maximum RKE-based diversity score. Additionally, we can consider a linear combination of RKE with the Precision, Density, or KID quality scores to identify a mixture of models that offers the best trade-off between quality and diversity.
Figure 2: Comparison of our proposed algorithms Mixture-UCB-CAB and Mixture-UCB-OGD with the baseline one arm online algorithms. We plot the diversity-measuring RKE score of the generated data and display 9 (randomly-selected) samples produced in the process of each algorithm.
@@
We perform several numerical experiments to test the application of our proposed Mixture-UCB approach in comparison to the Vanilla-UCB and One-Arm Oracle approaches that tend to generate samples from only one of the available generative models. Our numerical results indicate that the Mixture-UCB algorithms can generate samples with higher RKE diversity scores, and tends to generate samples from a mixture of several generative models when applied to image-based generative models. Also, we test the performance of Mixture-UCB on the KID, Precision, and Density scores, which similarly result in a higher score value for the mixture model found by the Mixture-UCB algorithm. We implement the Mixture-UCB by solving the convex optimization sub-problem at every iteration and also by applying the online gradient descent algorithm at every iteration. In our experiments, both implementations result in satisfactory results and can improve upon learning strategies tending to select only one generative model. The following is a summary of the contributions in this work:
@@
  • Studying the selection task for mixtures of multiple generative models to improve the diversity of generated samples.
@@
  • Proposing an online learning framework to address the mixture selection task for quadratic score functions.
@@
  • Developing the Mixture-UCB-CAB and Mixture-UCB-OGD algorithms to solve the formulated online learning problem.
  • Proving a finite horizon regret bound for Mixture-UCB-CAB which shows the convergence of Mixture-UCB-CAB to the optimal mixture.
  • Designing the Sparse-Mixture-UCB-CAB algorithm for the situation where we want to eventually select a small subset of the models. This is deferred to Appendix B.
@@
  • Presenting numerical results on the improvements in the diversity of generated data by the online selection of a mixture of the generation models.
@@ 2 Related Work
Assessment of Generative Models. The evaluation of generative models has been extensively studied, with a focus on both diversity and quality of generated images. Reference-free metrics such as Rényi Kernel Entropy (RKE) [7], VENDI [11], and the FKEA-VENDI [12] measure diversity without relying on ground-truth, while reference-based metrics such as Recall [9] and Coverage [10] assess diversity relative to real data. For image quality evaluation, Density and Precision metrics [10, 9] provide measures based on alignment with a reference distribution. The Wasserstein distance [13] and Fréchet Inception Distance (FID) [14] approximate the distance between real and generated datasets, while Kernel Inception Distance (KID) [6] uses squared maximum mean discrepancy for a kernel-based comparison of distributions. Also, the kernel-based KEN score [15] measures the novelty of generated samples produced by a generative model based on the entropy of their generated sample after canceling the mode-based diversity of a reference distribution.
@@
Multi-Armed Bandit Algorithms. The Multi-Armed Bandit (MAB) problem is a foundational topic in reinforcement learning, where an agent aims to maximize rewards from multiple options (arms) with initially unknown reward distributions [16, 17]. The Upper Confidence Bound (UCB) algorithm [18, 19, 20] is a widely adopted method for addressing the MAB problem, where uncertainty about an arm’s reward is replaced by an optimistic estimate. In generative models, optimism-based bandits have been applied to efficiently identify models with optimal Fréchet Inception Distance (FID) or Inception Score while minimizing data queries [5]. A special case of MAB, the continuum-armed bandit (CAB) problem [21], optimizes a function over continuous inputs, and has been applied to machine learning tasks such as hyperparameter optimization [22, 23]. Recent research explores CABs under more general smoothness conditions like Besov spaces [24], while other works have focused on regret bounds and Lipschitz conditions [25, 26, 27].
@@
Another related reference is informational multi-armed bandits [28], which extends UCB to maximizing the Shannon entropy of a discrete distribution, which is also a metric of diversity. In comparison, the algorithms in this paper can minimize the expectation of any quadratic positive semidefinite function, which not only covers the order-$2$ Rényi entropy for discrete distributions, but also includes the Rényi Kernel Entropy applicable to continuous data. Since the outputs of generative models are generally continuous, [28] is not applicable here. We also note that the recent paper [29] designs a contextual bandit algorithm for the online prompt-based selection of text-guided generative models, where the online learning algorithm adapts the selection of text-based generative models based on their UCB-estimated relevance CLIPScore in response to the current prompt. We note that our work studies a different online learning setting where the goal is to maximize the score of the mixture of unconditional generative models without any input text prompt.
@@ 3.1 Rényi Kernel Entropy
The Rényi Kernel Entropy [7] of the distribution $P$, which measures the diversity of the modes in $P$, is given by $\log(1/\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})])$, where $k$ is a positive definite kernel. [^2] The order-$2$ Rényi entropy for discrete distributions is a special case by taking $k(x,x^{\prime})=\mathbf{1}_{x=x^{\prime}}$. Taking the exponential of the Rényi Kernel Entropy, we have the RKE mode count $1/\mathbb{E}[k^{2}(X,X^{\prime})])$ [7], which is an estimate of the number of modes. Maximizing the RKE mode count is equivalent to minimizing the following loss
@@ 3.2 Maximum Mean Discrepancy and Kernel Inception Distance
The (squared) maximum mean discrepancy (MMD) [4] between distributions $P,Q$, which measures the distance between $P$ and $Q$, can be written as
@@
where $X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$ and $Y,Y^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}Q$, and $k$ is a positive definite kernel. Suppose $P$ is the distribution of samples from a generative model, and $Q$ is a reference distribution. Minimizing the MMD can ensure that $P$ is close to $Q$. The Kernel Inception Distance (KID) [6], a popular quality metric for image generative models, is obtained by first passing $P$ and $Q$ through the Inception network [30], and then computing their MMD, i.e., we have

### v2

# Be More Diverse than the Most Diverse: Optimal Mixtures of Generative Models via Mixture-UCB Bandit Algorithms

@@
Abstract
The availability of multiple training algorithms and architectures for generative models requires a selection mechanism to form a single model over a group of well-trained generation models. The selection task is commonly addressed by identifying the model that maximizes an evaluation score based on the diversity and quality of the generated data. However, such a best-model identification approach overlooks the possibility that a mixture of available models can outperform each individual model. In this work, we numerically show that a mixture of generative models on benchmark image datasets can indeed achieve a better evaluation score (based on FID and KID scores), compared to the individual models. This observation motivates the development of efficient algorithms for selecting the optimal mixture of the models. To address this, we formulate a quadratic optimization problem to find an optimal mixture model achieving the maximum of kernel-based evaluation scores including kernel inception distance (KID) and Rényi kernel entropy (RKE). To identify the optimal mixture of the models using the fewest possible sample queries, we view the selection task as a multi-armed bandit (MAB) problem and propose the Mixture Upper Confidence Bound (Mixture-UCB) algorithm that provably converges to the optimal mixture of the involved models. More broadly, the proposed Mixture-UCB can be extended to optimize every convex quadratic function of the mixture weights in a general MAB setting. We prove a regret bound for the Mixture-UCB algorithm and perform several numerical experiments to show the success of Mixture-UCB in finding the optimal mixture of text and image generative models. The project code is available at https://github.com/Rezaei-Parham/Mixture-UCB.
@@ 1 Introduction
The rapid advancements in generative modeling have created a need for mechanisms to combine multiple well-trained generative models, each developed using different algorithms and architectures, into a single unified model. Consider $m$ unconditional generative models $\mathcal{G}_{1},\ldots,\mathcal{G}_{m}$, where each $\mathcal{G}_{i}$ represents a probability model $P_{\mathcal{G}_{i}}$ according to which new samples are generated. A common approach for creating a unified model is to compute evaluation scores (e.g., the standard FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) scores) that quantify the diversity and fidelity of the generated data, followed by selecting the model $P_{\mathcal{G}_{i^{*}}}$ with the best evaluated score. This best-score model selection strategy has been widely adopted for choosing generative models across various domains, including image, text, and video data generation.
Figure 1: A mixture (right-most case) of FFHQ pre-trained generative models with weights (0.25,0.4,0.01,0.08,0.26) achieves better FID and KID scores compared to each of the five involved models. The mixture weights are computed using our proposed Mixture-UCB-OGD algorithm.
However, the model selection approach by identifying the score-maximizing model overlooks the possibility that a mixture of the generative models $\alpha_{1}P_{\mathcal{G}_{1}}+\cdots+\alpha_{m}P_{\mathcal{G}_{m}}$, where each sample is generated from a randomly-selected model with $\alpha_{i}$ being the probability of selecting model $\mathcal{G}_{i}$, can outperform every individual model. This motivates the following question: Can there be real-world settings where a non-degenerate mixture of some well-trained generative models obtain a better evaluation score compared to each individual model? Note that the standard FID and KID scores are convex functions of the generative model’s distribution, and thus they can be optimized by a non-degenerate mixture of the models. In this work, we numerically show that it is possible for a mixture of real-world generative models to improve evaluation scores over the individual models. An example is shown in Figure 1, where we find a non-degenerate mixture of five generative models pre-trained on the FFHQ dataset [^1] The pre-trained generative models are downloaded from dgm-eval GitHub repository (Stein et al., 2023).. As shown, assigning mixture weights $(0.25,0.4,0.01,0.08,0.26)$ to the models results in a significantly better FID score $170.11$ than the best individual FID $185.07$.
To understand the improvement achieved by the mixture model, we note that the FID and KID scores evaluate both the quality and diversity of the generated data. While the averaged quality of generated samples represents an expected value that is optimized by an individual model, the diversity of samples from a mixture of the models can significantly improve over the diversity of the individual models’ data. Figure 2 displays an illustrative example for this point, where we observe that the diversity of “red bird, cartoon style” samples generated by each of the three text-to-image models, is qualitatively and quantitatively [^2] We have evaluated the diversity scores RKE (Jalali et al., 2023) and Vendi (Dan Friedman & Dieng, 2023). lower than the diversity of their mixture. As a result, the improvement in the diversity of a mixture of generative models can result in an improved FID and KID evaluation scores, as we numerically observe in Figure 1.
1.1 Computing Optimal Mixtures of Generative Models: The Mixture-UCB Multi-Armed Bandit Algorithm
Since the evaluation score of a mixture of generative models can improve over the scores of the individual models, a natural question is how to efficiently compute the weights of an optimal mixture of the models using the fewest possible samples from the models. Here, our goal is to minimize the number of sample generation queries from sub-optimal models, which will save the time and monetary costs of identifying the best model. To achieve this, we propose viewing the task as a multi-armed bandit (MAB) problem, in which every generative model represents an arm and our goal is to find the best mixture of the models with the optimal evaluation score. The MAB approach for selecting among generative models has been recently explored by Hu et al. (2024) applying the Upper Confidence Bound (UCB) algorithm to the FID score. This MAB-based model selection extends the online model selection methods for supervised models, including the successive halving strategy (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) [^3] Jamieson & Talwalkar (2016) focused on applying successive halving on hyperparameter optimization for supervised learning, whereas Chen & Ghosh (2024) focused on generative models using maximum mean discrepancy (Gretton et al., 2012) as the score..
@@ 1.1 Computing Optimal Mixtures of Generative Models: The Mixture-UCB Multi-Armed Bandit Algorithm
However, in the existing MAB algorithms, the goal is to eventually converge to a single arm with the optimal score. Successive halving (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) and the UCB algorithm developed by Hu et al. (2024) will ultimately select only one generative model after a sufficient number of iterations. However, as we discussed earlier, the evaluation scores of the generated data could be higher when the sample generation follows a mixture of models rather than a single model. This observation leads to the following task: Developing an MAB algorithm that finds the optimal mixture of the arms rather than the single best arm.
@@
In this work, we develop a general MAB method, which we call the Mixture-UCB algorithm, that can provably find the best mixture of the arms when the evaluation score is a quadratic function of the arms’ distribution, i.e. when it represents the average of scores assigned to the pairs of drawn samples. Formulating the optimization problem for a quadratic score function results in a quadratic online convex optimization problem that can be efficiently solved using the online gradient descent algorithm. More importantly, we establish a concentration bound for the quadratic function of the mixture weights, which enables us to extend the UCB algorithm to the Mixture-UCB method for the online selection of the mixture weights.
@@
For selecting mixtures of generative models using Mixture-UCB, we focus on evaluation scores that reduce to a quadratic function of the generative model’s distribution, including Kernel Inception Distance (KID) (Bińkowski et al., 2018), Rényi Kernel Entropy (RKE) (Jalali et al., 2023) scores, as well as the quality-measuring Precision (Sajjadi et al., 2018; Kynkäänniemi et al., 2019) and Density (Naeem et al., 2020) scores, which are linear functions of the generative distribution. Among these scores, RKE provides a reference-free entropy function for assessing the diversity of generated data, making it suitable for quantifying the variety of generated samples. Our mixture-based MAB framework can therefore be applied to find the mixture model with the maximum RKE-based diversity score. Additionally, we consider a linear combination of RKE with the Precision, Density, or KID quality scores to find a mixture of models that offers the best trade-off between quality and diversity.
Figure 2: Visual comparison of the diversity across individual arms and the optimal mixture for images generated using models Kandinsky 3, Stable Diffusion 3, and PixArt-$\alpha$ with the prompt “Red bird, cartoon style”. The mixture weights are computed via the Mixture-UCB-OGD method.
@@
We perform several numerical experiments to test the application of our proposed Mixture-UCB approach in comparison to the Vanilla-UCB and One-Arm Oracle approaches that tend to generate samples from only one of the available generative models. Our numerical results indicate that the Mixture-UCB algorithms can generate samples with higher RKE diversity scores, and tends to generate samples from a mixture of several generative models when applied to image-based generative models. Also, we test the performance of Mixture-UCB on the KID, Precision, and Density scores, which similarly result in a higher score value for the mixture model found by the Mixture-UCB algorithm. We implement the Mixture-UCB by solving the convex optimization sub-problem at every iteration and also by applying the online gradient descent algorithm at every iteration. In our experiments, both implementations result in satisfactory results and can improve upon learning strategies tending to select only one generative model. Here is a summary of this work’s contributions:
@@
  • Studying the selection task for mixtures of multiple generative models to improve the evaluation scores of generated samples (Section 4).
@@
  • Proposing an online learning multi-armed bandit framework to address the mixture selection task for quadratic score functions (Section 5).
@@
  • Developing the Mixture-UCB-CAB and Mixture-UCB-OGD algorithms to solve the formulated online learning problem and proving a regret bound for Mixture-UCB-CAB (Sections 5.1, 5.2).
@@
  • Presenting numerical results on the improvements in the diversity of generated data by the online selection of a mixture of the generation models (Section 6, Appendix 8.4).
@@ 2 Related Work
Assessment of Generative Models. The evaluation of generative models has been extensively studied, with a focus on both diversity and quality of generated images. Reference-free metrics such as Rényi Kernel Entropy (RKE) (Jalali et al., 2023) and VENDI (Dan Friedman & Dieng, 2023; Ospanov et al., 2024) measure diversity without relying on ground-truth, while reference-based metrics such as Recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019) and Coverage (Naeem et al., 2020) assess diversity relative to real data. For image quality evaluation, Density and Precision metrics (Naeem et al., 2020; Kynkäänniemi et al., 2019) provide measures based on alignment with a reference distribution. The Wasserstein distance (Arjovsky et al., 2017) and Fréchet Inception Distance (FID) (Heusel et al., 2017) approximate the distance between real and generated datasets, while Kernel Inception Distance (KID) (Bińkowski et al., 2018) uses squared maximum mean discrepancy for a kernel-based comparison of distributions. Wang et al. (2023) apply the KID score for the distributed evaluation of generative models. The novelty evaluation of generative models has also been studied in References (Han et al., 2022; Jiralerspong et al., 2023; Zhang et al., 2024b; a).
@@
Multi-Armed Bandit Algorithms. The Multi-Armed Bandit (MAB) problem is a foundational topic in reinforcement learning, where an agent aims to maximize rewards from multiple options (arms) with initially unknown reward distributions (Lai & Robbins, 1985; Thompson, 1933). The Upper Confidence Bound (UCB) algorithm (Agrawal, 1995a; Auer, 2002; Bubeck et al., 2012) is a widely adopted method for addressing the MAB problem, where uncertainty about an arm’s reward is replaced by an optimistic estimate. In generative models, optimism-based bandits have been applied to efficiently identify models with optimal Fréchet Inception Distance (FID) or Inception Score while minimizing data queries (Hu et al., 2024). A special case of MAB, the continuum-armed bandit (CAB) problem (Agrawal, 1995b), optimizes a function over continuous inputs, and has been applied to machine learning tasks such as hyperparameter optimization (Feurer & Hutter, 2019; Li et al., 2018). Recent research explores CABs under more general smoothness conditions like Besov spaces (Singh, 2021), while other works have focused on regret bounds and Lipschitz conditions (Kleinberg, 2004; Kleinberg et al., 2019; Bubeck et al., 2008).
@@
Another related reference is informational multi-armed bandits (Weinberger & Yemini, 2023), which extends UCB to maximizing the Shannon entropy of a discrete distribution. In comparison, the algorithms in this paper can minimize the expectation of any quadratic positive-semidefinite function, which also covers the order-$2$ Rényi entropy for discrete distributions. Since the generative models’ outputs are generally continuous, (Weinberger & Yemini, 2023) is not applicable to our setting.
@@ 3.1 Rényi Kernel Entropy
The Rényi Kernel Entropy (Jalali et al., 2023) of the distribution $P$, which measures the diversity of the modes in $P$, is given by $\log(1/\mathbb{E}_{X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P}[k^{2}(X,X^{\prime})])$, where $k$ is a positive definite kernel. [^4] The order-$2$ Rényi entropy for discrete distributions is a special case by taking $k(x,x^{\prime})=\mathbf{1}_{x=x^{\prime}}$. Taking the exponential of the Rényi Kernel Entropy, we have the RKE mode count $1/\mathbb{E}[k^{2}(X,X^{\prime})])$ (Jalali et al., 2023), which is an estimate of the number of modes. Maximizing the RKE mode count is equivalent to minimizing the following loss
@@ 3.2 Maximum Mean Discrepancy and Kernel Inception Distance
The (squared) maximum mean discrepancy (MMD) (Gretton et al., 2012) between distributions $P,Q$, which measures the distance between $P$ and $Q$, can be written as
@@
where $X,X^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P$ and $Y,Y^{\prime}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}Q$, and $k$ is a positive definite kernel. Suppose $P$ is the distribution of samples from a generative model, and $Q$ is a reference distribution. Minimizing the MMD can ensure that $P$ is close to $Q$. The Kernel Inception Distance (KID) (Bińkowski et al., 2018), a popular quality metric for image generative models, is obtained by first passing $P$ and $Q$ through the Inception network (Szegedy et al., 2016), and then computing their MMD, i.e., we have